# Fast Multiplication and the PLWE–RLWE Equivalence for an Infinite Family of Cyclotomic Subextensions

### v1

@@ 1 Introduction
Out of the first three post-quantum cryptography (PQC) standards published in 2024 by the National Institute of Standards and Technologies (NIST), two of them, namely ML-KEM [FIPS:2024:203] and ML-DSA [FIPS:2024:204], belong to the lattice-based family. This is not surprising, as this family has been dominant throughout all the stages of the standardization process: lattice-based cryptography has been the paradigm which has had the most schemes based upon, up until the final rounds. This is due to the fact that lattice-based post-quantum primitives can provide both acceptable key sizes and efficient performance while maintaining a solid understanding of the underlying mathematical hardness. The fact that lattice-based schemes enjoy homomorphic properties and can be turned into Fully Homomorphic Encryption (FHE) primitives is another important reason for the prevalence of the lattice family.
@@ 1.1 LWE and its structured variants
One should note that for Galois extensions, the search problem can be effectively reduced to the decision problem [EHL]. An even stronger result exists [PRSD:2017:GRLWE], providing a reduction for any ring and modulus. Let us also mention that, for the RLWE distribution, we have given the definition of the primal version. Another definition can be considered, by taking the secret $s$ and the support of $\chi$ over the trace-dual of the ring of integers $\mathcal{O}_{K}^{\vee}$. This is indeed the original definition of the RLWE problem in [LPR:2013:ILL]. However, both problems were shown to be equivalent in [RSW:2018:RPP], and hence we will stick to the primal version for the sake of simplicity.
@@
It is worth noting that this cryptographic family has been found extremely useful not only in PQC but also in other relevant aspects of cryptography, most notably in Homomorphic Encryption (HE). This type of cryptography inherits its name from the mathematical notion of homomorphisms: intuitively, it allows to perform certain operations on the encrypted plaintext, in the same way as if we decrypt the ciphertext and apply the operation over the plaintext, mimicking the properties of homomorphic maps. The impact of practical schemes that achieve this property cannot be overstated, especially in fields in which the privacy of data is of the utmost importance. Depending on the nature of the operations allowed, we can distinguish between partially homomorphic encryption or fully homomorphic encryption, among others. The first provably lattice-based homomorphic encryption scheme is the GPV scheme [10.1145/1374376.1374407]. Thereafter, a vast number of HE schemes have been derived from the LWE paradigm and its structured variants [10.1145/2090236.2090262, BFV:2012:144].
@@
The structure added in these variants gives us the benefit of reductions to approximate lattice problems like the shortest vector problem (SVP) over ideal lattices. However, the added algebraic structure can be a source of additional attacks that are ineffective against the unstructured variants. Particularly for PLWE, there exists a number of relevant attacks which employ information about the roots of the polynomial $f(x)$ behind the PLWE instance. In [ELOS:2015:PWI, ELOS:2016:RCN], a number of attacks against the decisional (and sometimes the search) version of PLWE are presented, if the polynomial $f(x)$ has a root $\alpha$ of small order or of small residue in $\mathbb{F}_{q}$, by constructing a number of distinguishability sets made from the evaluation of the samples on the critical root. Moreover, a more subtle family of statistical attacks can be deployed under certain conditions discussed therein. More recently, in [BDM:2024:GARB], these attacks (those based on roots of small order/residue as well as the statistical ones) have been generalized to polynomials with roots belonging to arbitrary degree field extensions of $\mathbb{F}_{q}$.
@@ 1.2 Relation between RLWE and PLWE problems
The RLWE and PLWE paradigms were introduced roughly at the same time. In 2009, [SSTX:2009:EPK] presented a structured variant of the LWE problem based on ideal lattices over rings of the form $\mathcal{O}=\mathbb{Z}[x]/(f(x))$. They focused on cyclotomic polynomials giving rise to the first technical description of the PLWE paradigm.
Meanwhile, in 2010, [LPR:2013:ILL] introduced another structured variant of LWE, this time over the ring of integers of algebraic number fields. While most of the results in this work were generalized to arbitrary number fields, the complete reduction involving the pseudorandomness of the RLWE distribution, an important feat with regards to practical cryptographic instances, was only proved for cyclotomic fields. The complete arbitrary number field generalization that lifted the requirement of cyclotomic fields was proved in [PRSD:2017:GRLWE].
As noted above, both approaches have important cryptographic qualities: the RLWE problem bases its security on a more established and studied problem and has not been prone to many attacks over its structure. This is contrary to existing approaches towards exploiting the information given in the PLWE problem, such as root-based approaches [ELOS:2015:PWI, ELOS:2016:RCN, BDM:2024:GARB].
@@
The existence of equivalence between RLWE and PLWE for different families of number fields has been an important object of study. In [RSW:2018:RPP, Section 4], this equivalence was proved for two reduced but large families of polynomials. For cyclotomic polynomials, [SSS:2024:RPNE] showed that when no restriction is imposed on the conductor of the cyclotomic field, the two problems are not equivalent in general. However, if certain conditions are imposed on the conductor, it can be shown that these problems are indeed equivalent. In [Blanco:2020:REC], the author shows that if the conductor is divisible by a bounded number of primes, the condition number of the associated powers-of-the-roots Vandermonde matrix which reigns over the equivalence (which represents the variation induced from the difference between the coordinate and canonical embedding associated with the PLWE and RLWE problems, respectively) is bounded by a polynomial factor and, therefore, the problems are equivalent.
Another topic which has recently gained traction is the study of the R/PLWE problem for the maximal totally real subextension of a cyclotomic field. The interest for this field was introduced in [BL:2022:RPE, Blanco:2021:VMTR], where it was proved that the defining polynomials for these subfields are resistant to PLWE root-based attacks for $\alpha=\pm 2$ as well as for $\alpha=\pm 1$.
Furthermore, it is a strategic approach to have at hand as large a family of fields and rings as possible, with the idea of being prepared for a potential catastrophic attack against current structured cryptosystems. Indeed, it is worth to mention [cramerducaswesolowsky], where the authors exploit certain class-field-theoretical properties of the Stickelberger ideal to prove the following fact:
@@
$||v||\leq N\mathfrak{a}^{1/n}\exp(\mathcal{O}(\sqrt{n}).$
Even more striking is the recent refinement of Theorem 1.1, which shows that under mild but technical conditions, the quantum algorithm runs in probabilistic polynomial time in $[K:\mathbb{Q}]$, $h_{K}^{+}$ and $\log(N\mathfrak{a})$, where $h_{K}^{+}$ is the narrow class number of $K$. The reader is referred to [CDW2, Theorem 5.1] for further details.
@@ 1.3 Contributions and Organization
Previous works related to ours include [Blanco:2021:VMTR] and [BL:2022:RPE] addressing the RLWE–PLWE equivalence for maximal totally real cyclotomic subextensions, and [lyubashevsky2019nttru], where an NTT-based multiplication scheme is proposed for the non-power-of-two $2^{r}3^{s}$-th cyclotomic fields, with $r,s\geq 1$.
@@ 2 The Discrete Cosine Transform, $\mathsf{DCT}$
The Discrete Cosine Transform ($\mathsf{DCT}$) is widely used in many digital signal processing applications. Multiple fast algorithms for different types of the Discrete Cosine Transform have been reported in the literature [bi1998dct, hou1987fast, kok1997fast, ahmed1974discrete]. In this paper, we shall use types II and III, which are mutual inverses when scaled properly.
@@
The inverse of the type-III $\mathsf{DCT}$ is given by the type-II $\mathsf{DCT}$. The non-scaled type-II Discrete Cosine Transformation of the sequence $a(k)$ is a new sequence $a^{\prime}(j)$ given by
@@ 3 RLWE–PLWE equivalence for $n=2^{r}3^{s}$
The study of the equivalence between RLWE and PLWE was first addressed in [DD:2017] for cyclotomic number fields of degree $2^{k}p$ or $2^{k}pq$ with $p,q$ primes and $q<p$, and later on in [RSW:2018:RPP, Section 4] for two particular non-cyclotomic families of polynomials. In [DSS:2022], the authors obtained an explicit closed formula for the condition number of the $2^{k}p^{l}$ cyclotomic field, and in [SSS:2024:RPNE] the authors showed that for an arbitrary conductor, the two problems are not equivalent for cyclotomic fields.
Still within the cyclotomic family, in [Blanco:2020:REC], it was proved that if the conductor is divisible by a bounded number of primes, the RLWE–PLWE equivalence holds alongside with much sharper bounds for the condition number for conductors divisible by up to $3$ primes. This result was recently generalized in [BPBN] to conductors divisible by up to $6$ primes. The authors also introduce and study cyclo-multiquadratic number fields, a family which also enjoys RLWE–PLWE equivalence and fast multiplication.
Beyond cyclotomics, in [Blanco:2021:VMTR], the RLWE–PLWE equivalence was proved for the maximal totally real subfield of the $4p$-th cyclotomic field. This result was generalized in [BL:2022:RPE] to conductors of the form $2^{r}pq$ with $p<q$ primes, as well as for $p=1$ or $q=1$. Furthermore, in this last work it was proved that the defining polynomials of these subfields are resistant to the PLWE root-based attacks for $\alpha=\pm 1,\pm 2$. This gives a reason to pay attention to these fields from a cryptographic point of view. However, cyclotomic polynomials can vanish at $\alpha=2$ modulo suitable primes, and the authors provided some examples.
To set the stage for our result, let $\psi_{n}=\zeta_{n}+\zeta_{n}^{-1}=2\cos(2\pi/n)$ so $\mathbb{Q}(\psi_{n})=\mathbb{Q}(\zeta_{n})^{+}$ is the maximal real subfield of the cyclotomic field $\mathbb{Q}(\zeta_{n})$. We know that $[\mathbb{Q}(\psi_{n}):\mathbb{Q}]=\phi(n)/2$, where $\phi$ is Euler’s totient function. It is well known that the ring of integers of $\mathbb{Q}(\psi_{n})$ is $\mathbb{Z}[\psi_{n}]\cong\mathbb{Z}[x]/(\Psi_{n}(x))$, where $\Psi_{n}(x)$ is the minimal polynomial of $\psi_{n}$ of degree $\phi(n)/2$. For these standard claims, we refer to [washington2012introduction].
@@
To show that the PLWE and RLWE problems are equivalent, it enough to prove that the condition number of the matrix $\mathbf{M}$ is bounded by a polynomial in $n$. In this case, we say that $\mathbf{M}$ is well conditioned. To see this, first note that the matrix $\mathbf{M}$ is well conditioned if and only if the matrix
@@ 4 Fast multiplication in $\mathbb{Z}[x]/(\Psi_{n}(x))$ via the $\mathsf{DCT}$
In this section, we introduce an algorithm for fast multiplication over the ring $\mathbb{Z}[x]/(\Psi_{n}(x))$. The algorithm is similar to the Number Theoretic Transform (NTT). However, the NTT is typically computed over the ring $\mathbb{Z}[x]/(x^{n}+1)$, where $n$ is a power of two yielding the ring of integers of the $2n$-th cyclotomic field [agarwal1975ntt, pollard1971fast]. For non-power-of-two cyclotomic fields, in [lyubashevsky2019nttru] the authors describe an NTT-based multiplication algorithm in the $2^{r}3^{s}$-th cyclotomic field, where $r\geq 1$ and $s\geq 1$. The algorithms developed to compute the $\mathsf{DCT}$ and its inverse $\mathsf{IDCT}$ can be classified into two categories: indirect and direct. Indirect algorithms take advantage of existing fast algorithms, such as the fast Fourier and Hadamard transforms [ahmed1974discrete]. On the other hand, direct algorithms reduces the computational complexity by means of matrix factorization and recursive decomposition as a generalization of the Cooley–Tukey algorithms with the same asymptotic complexity $\mathcal{O}(n\log n)$ [bi1998dct, bi1999fast, hou1987fast, kok1997fast].
@@ 4.1 Case $n=2^{r}$
It is well known that $\Psi_{n}(x)=V_{m}(x)$ [loper2016resultants, Theorem 2.6] where $V_{m}(x)$ is the modified Chebyshev polynomial from Definition 3.1. Every element in the ring $R:=\mathbb{Z}[x]/(\Psi_{n}(x))$ can be represented by a polynomial of degree less than or equal to $m-1$. Furthermore, if $p(x),q(x)\in R$ and $N=m$, then $\mathsf{DCT}(p(x))$ and $\mathsf{DCT}(q(x))$ can be computed in order $n\log n$ [bi1998dct, bi1999fast].
@@
The fast implementation of the inverse transform $\mathsf{IDCT}$ requires in total $(r-2)2^{r-2}$ multiplications and $13(r-2)2^{r-3}-2^{r-2}+$ 1 additions [kok1997fast, hou1987fast]. This means the asymptotic complexity of computing $\mathbf{c}$ is $\mathcal{O}(m\log m)$.

### v2

@@ Abstract
keywords:
Number Theoretic Transform, Discrete Cosine Transform, Fast multiplication, Ring Learning with Errors, Polynomial Learning with Errors, Condition number, Abelian number fields
Acknowledgements
I. Blanco-Chacón, W. Bolaños, A. Haavikko, and C. Hollanti were supported in part by the Research Council of Finland (grant #351271). C. Hollanti was also supported by the Finnish Ministry of Defence, MATINE (grant #2500M-0147). I. Blanco-Chacón would like to acknowledge the visiting professor funding from the Aalto Science Institute and the Department of Mathematics and Systems Analysis at Aalto University, Finland. R. M. Sánchez-Ledesma was supported in part by the PQReact project, which has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement #101119547.
Statements and Declarations
There are no competing financial nor non-financial interests that are directly or indirectly related to the work submitted for publication.
@@ 1 Introduction
Out of the first three post-quantum cryptography (PQC) standards published in 2024 by the National Institute of Standards and Technologies (NIST), two of them, namely ML-KEM [1] and ML-DSA [2], belong to the lattice-based family. This is not surprising, as this family has been dominant throughout all the stages of the standardization process: lattice-based cryptography has been the paradigm which has had the most schemes based upon, up until the final rounds. This is due to the fact that lattice-based post-quantum primitives can provide both acceptable key sizes and efficient performance while maintaining a solid understanding of the underlying mathematical hardness. The fact that lattice-based schemes enjoy homomorphic properties and can be turned into Fully Homomorphic Encryption (FHE) primitives is another important reason for the prevalence of the lattice family.
@@ 1.1 LWE and its structured variants
In literature, the most common choice is to use cyclotomic extensions $K=\mathbb{Q}(\zeta_{2n})$ with minimal polynomials $f(x)=x^{n}+1$, where $n$ is a power of two or similarly the polynomial quotient rings $\mathcal{O}_{f}=\mathbb{Z}[x]/(x^{n}+1)$. In fact, for this choice of $K$ and $f(x)$, the two rings $\mathcal{O}_{K}$ and $\mathcal{O}_{f}$ are isomorphic.
Note that for Galois extensions, the search problem can be effectively reduced to the decision problem [3]. An even stronger result exists [4], providing a reduction for any ring and modulus. Let us also mention that, for the RLWE distribution, we have given the definition of the primal version. Another definition can be considered, by taking the secret $s$ and the support of $\chi$ over the trace-dual of the ring of integers $\mathcal{O}_{K}^{\vee}$. This is indeed the original definition of the RLWE problem in [5]. However, both problems were shown to be equivalent in [6], and hence we will stick to the primal version for the sake of simplicity.
@@
It is worth noting that this cryptographic family has been found extremely useful not only in PQC but also in other relevant aspects of cryptography, most notably in Homomorphic Encryption (HE). This type of cryptography inherits its name from the mathematical notion of homomorphisms: intuitively, it allows to perform certain operations on the encrypted plaintext, in the same way as if we decrypt the ciphertext and apply the operation over the plaintext, mimicking the properties of homomorphic maps. The impact of practical schemes that achieve this property cannot be overstated, especially in fields in which the privacy of data is of the utmost importance. Depending on the nature of the operations allowed, we can distinguish between partially homomorphic encryption or fully homomorphic encryption, among others. The first provably lattice-based homomorphic encryption scheme is the GPV scheme [7]. Thereafter, a vast number of HE schemes have been derived from the LWE paradigm and its structured variants [8, 9, 10, 11, 12, 13].
@@
The structure added in these variants gives us the benefit of reductions to approximate lattice problems like the shortest vector problem (SVP) over ideal lattices. However, the added algebraic structure can be a source of additional attacks that are ineffective against the unstructured variants. Particularly for PLWE, there exists a number of relevant attacks which employ information about the roots of the polynomial $f(x)$ behind the PLWE instance. In [14, 15], a number of attacks against the decisional (and sometimes the search) version of PLWE are presented, if the polynomial $f(x)$ has a root $\alpha$ of small order or of small residue in $\mathbb{F}_{q}$, by constructing a number of distinguishability sets made from the evaluation of the samples on the critical root. Moreover, a more subtle family of statistical attacks can be deployed under certain conditions discussed therein. More recently, in [16], these attacks (those based on roots of small order/residue as well as the statistical ones) have been generalized to polynomials with roots belonging to arbitrary degree field extensions of $\mathbb{F}_{q}$.
@@ 1.2 Relation between RLWE and PLWE problems
The RLWE and PLWE paradigms were introduced roughly at the same time. In 2009, [17] presented a structured variant of the LWE problem based on ideal lattices over rings of the form $\mathcal{O}=\mathbb{Z}[x]/(f(x))$. They focused on cyclotomic polynomials giving rise to the first technical description of the PLWE paradigm.
Meanwhile, in 2010, [5] introduced another structured variant of LWE, this time over the ring of integers of algebraic number fields. While most of the results in this work were generalized to arbitrary number fields, the complete reduction involving the pseudorandomness of the RLWE distribution, an important feat with regards to practical cryptographic instances, was only proved for cyclotomic fields. The complete arbitrary number field generalization that lifted the requirement of cyclotomic fields was proved in [4].
As noted above, both approaches have important cryptographic qualities: the RLWE problem bases its security on a more established and studied problem and has not been prone to many attacks over its structure. This is contrary to existing approaches towards exploiting the information given in the PLWE problem, such as root-based approaches [14, 15, 16].
@@
The existence of equivalence between RLWE and PLWE for different families of number fields has been an important object of study. In [6, Section 4], this equivalence was proved for two reduced but large families of polynomials. For cyclotomic polynomials, [18] showed that when no restriction is imposed on the conductor of the cyclotomic field, the two problems are not equivalent in general. However, if certain conditions are imposed on the conductor, it can be shown that these problems are indeed equivalent. In [19], the author shows that if the conductor is divisible by a bounded number of primes, the condition number of the associated powers-of-the-roots Vandermonde matrix which reigns over the equivalence (which represents the variation induced from the difference between the coordinate and canonical embedding associated with the PLWE and RLWE problems, respectively) is bounded by a polynomial factor and, therefore, the problems are equivalent.
Another topic which has recently gained traction is the study of the R/PLWE problem for the maximal totally real subextension of a cyclotomic field. The interest for this field was introduced in [20, 21], where it was proved that the defining polynomials for these subfields are resistant to PLWE root-based attacks for $\alpha=\pm 2$ as well as for $\alpha=\pm 1$.
Furthermore, it is a strategic approach to have at hand as large a family of fields and rings as possible, with the idea of being prepared for a potential catastrophic attack against current structured cryptosystems. Indeed, it is worth to mention [22], where the authors exploit certain class-field-theoretical properties of the Stickelberger ideal to prove the following fact:
@@
$||v||\leq N\mathfrak{a}^{1/n}\exp(\mathcal{O}(\sqrt{n})).$
Even more striking is the recent refinement of Theorem 1.1, which shows that under mild but technical conditions, the quantum algorithm runs in probabilistic polynomial time in $[K:\mathbb{Q}]$, $h_{K}^{+}$ and $\log(N\mathfrak{a})$, where $h_{K}^{+}$ is the narrow class number of $K$. The reader is referred to [23, Theorem 5.1] for further details.
@@ 1.3 Contributions and Organization
Previous works related to ours include [21] and [20] addressing the RLWE–PLWE equivalence for maximal totally real cyclotomic subextensions, and [24], where an NTT-based multiplication scheme is proposed for the non-power-of-two $2^{r}3^{s}$-th cyclotomic fields, with $r,s\geq 1$.
@@ 2 The Discrete Cosine Transform, $\mathsf{DCT}$
The Discrete Cosine Transform ($\mathsf{DCT}$) is widely used in many digital signal processing applications. Multiple fast algorithms for different types of the Discrete Cosine Transform have been reported in the literature [25, 26, 27, 28]. In this paper, we shall use types II and III, which are mutual inverses when scaled properly.
@@
The inverse of the type-III $\mathsf{DCT}$ is given by the type-II $\mathsf{DCT}$. The non-scaled type-II Discrete Cosine Transform of the sequence $a(k)$ is a new sequence $a^{\prime}(j)$ given by
@@ 3 RLWE–PLWE equivalence for $n=2^{r}3^{s}$
The study of the equivalence between RLWE and PLWE was first addressed in [29] for cyclotomic number fields of degree $2^{k}p$ or $2^{k}pq$ with $p,q$ primes and $q<p$, and later on in [6, Section 4] for two particular non-cyclotomic families of polynomials. In [30], the authors obtained an explicit closed formula for the condition number of the $2^{k}p^{l}$ cyclotomic field, and in [18] the authors showed that for an arbitrary conductor, the two problems are not equivalent for cyclotomic fields.
Still within the cyclotomic family, in [19], it was proved that if the conductor is divisible by a bounded number of primes, the RLWE–PLWE equivalence holds alongside with much sharper bounds for the condition number for conductors divisible by up to $3$ primes. This result was recently generalized in [31] to conductors divisible by up to $6$ primes. The authors also introduce and study cyclo-multiquadratic number fields, a family which also enjoys RLWE–PLWE equivalence and fast multiplication.
Beyond cyclotomics, in [21], the RLWE–PLWE equivalence was proved for the maximal totally real subfield of the $4p$-th cyclotomic field. This result was generalized in [20] to conductors of the form $2^{r}pq$ with $p<q$ primes, as well as for $p=1$ or $q=1$. Furthermore, in this last work it was proved that the defining polynomials of these subfields are resistant to the PLWE root-based attacks for $\alpha=\pm 1,\pm 2$. This gives a reason to pay attention to these fields from a cryptographic point of view. However, cyclotomic polynomials can vanish at $\alpha=2$ modulo suitable primes, and the authors provided some examples.
To set the stage for our result, let $\psi_{n}=\zeta_{n}+\zeta_{n}^{-1}=2\cos(2\pi/n)$ so $\mathbb{Q}(\psi_{n})=\mathbb{Q}(\zeta_{n})^{+}$ is the maximal real subfield of the cyclotomic field $\mathbb{Q}(\zeta_{n})$. We know that $[\mathbb{Q}(\psi_{n}):\mathbb{Q}]=\phi(n)/2$, where $\phi$ is Euler’s totient function. It is well known that the ring of integers of $\mathbb{Q}(\psi_{n})$ is $\mathbb{Z}[\psi_{n}]\cong\mathbb{Z}[x]/(\Psi_{n}(x))$, where $\Psi_{n}(x)$ is the minimal polynomial of $\psi_{n}$ of degree $\phi(n)/2$. For these standard claims, we refer to [32].
@@
To show that the PLWE and RLWE problems are equivalent, it is enough to prove that the condition number of the matrix $\mathbf{M}$ is bounded by a polynomial in $n$. We adapt this approach from [29, 6]. For a short summary, see [21] and the discussion following Definition 2.3. If we can find a polynomial bound in $n$, we say that $\mathbf{M}$ is well conditioned. To see that $\mathbf{M}$ is well conditioned, first note that the matrix $\mathbf{M}$ is well conditioned if and only if the matrix
@@ 4 Fast multiplication in $\mathbb{Z}[x]/(\Psi_{n}(x))$ via the $\mathsf{DCT}$
In this section, we introduce an algorithm for fast multiplication over the ring $\mathbb{Z}[x]/(\Psi_{n}(x))$. The algorithm is similar to the Number Theoretic Transform (NTT). However, the NTT is typically computed over the ring $\mathbb{Z}[x]/(x^{n}+1)$, where $n$ is a power of two yielding the ring of integers of the $2n$-th cyclotomic field [33, 34]. For non-power-of-two cyclotomic fields, in [24] the authors describe an NTT-based multiplication algorithm in the $2^{r}3^{s}$-th cyclotomic field, where $r\geq 1$ and $s\geq 1$. The algorithms developed to compute the $\mathsf{DCT}$ and its inverse $\mathsf{IDCT}$ can be classified into two categories: indirect and direct. Indirect algorithms take advantage of existing fast algorithms, such as the fast Fourier and Hadamard transforms [28]. On the other hand, direct algorithms reduce the computational complexity by means of matrix factorization and recursive decomposition as a generalization of the Cooley–Tukey algorithm with the same asymptotic complexity $\mathcal{O}(n\log n)$ [25, 35, 26, 27].
@@ 4.1 Case $n=2^{r}$
It is well known that $\Psi_{n}(x)=V_{m}(x)$ [36, Theorem 2.6] where $V_{m}(x)$ is the modified Chebyshev polynomial from Definition 3.1. Every element in the ring $R:=\mathbb{Z}[x]/(\Psi_{n}(x))$ can be represented by a polynomial of degree less than or equal to $m-1$. Furthermore, if $p(x),q(x)\in R$ and $N=m$, then $\mathsf{DCT}(p(x))$ and $\mathsf{DCT}(q(x))$ can be computed in order $n\log n$ [25, 35].
@@
The fast implementation of the inverse transform $\mathsf{IDCT}$ requires in total $(r-2)2^{r-2}$ multiplications and $13(r-2)2^{r-3}-2^{r-2}+$ 1 additions [27, 26]. This means the asymptotic complexity of computing $\mathbf{c}$ is $\mathcal{O}(m\log m)$.